OpenStudy (anonymous):

There are 15 employees that must attend training sessions where only 3 people can be trained at a time. How many different possible groups are there?

OpenStudy (anonymous):

15/3= 5

OpenStudy (anonymous):

1 x 15 = 15 psossible groups (pg) 3 x 5 = 5 pg 2 x 7 + 1 = 8 pg 2 X 6 + 3 = 7 pg 2 x 5 + 3 + 2 = 7 pg realy it goes on for a long time

OpenStudy (anonymous):

so how do i find the answer

OpenStudy (anonymous):

This is a factorial problem. You have three spaces, you have 15 people to choose from for the first space and 14 people to choose from in the second space and 13 people to choose from for the third space.

OpenStudy (anonymous):

Also, realize that a group of people #1, #2, and #3 is the same as a group of #3, #2, and #1. So, you need to divide by the possible number of duplicates.